HUBBLE SPACE TELESCOPE DAILY REPORT #1695 PERIOD COVERED: 0000Z (UTC) 08/12/96 - 0000Z (UTC) 08/13/96 Daily Status Report as of 226/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Competed Two sets of FOS/RD 6552 (Spectrophotometry of Inner = Jovian Satellites) The Faint Object Spectrograph was used to make spectrographic = observations of Jupiter's inner moon Thebe during its eastern elongation. = Many things can be determined from the analysis of these spectra. The = observations were executed as planned, and no problems were reported. 1.2 Completed GHRS/2 6541 (Interstellar Cadmium: Probing Galactic = Chemical Evolution) The High Resolution Spectrograph (Side Two) was used to make spectr= ographic observations of the O/B class star HD144217 in order to probe = interstellar Cadmium and study the galactic chemical evolution. The = observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6353 (A Snapshot Emission Line Imaging Survey of = Very Low Excitation Planetary Nebulas) The WF/PC-2 was used to make snapshot observations as part to an = emission line survey of the very low excitation planetary nebula = PK327-02D1 (Class PNe). This target was selected based on the O III/H = alpha line flux ratio. The observations were executed as scheduled, and = no anomalies were noted. 1.4 Completed FOS/BL 6498 (Accretion Streams in Magnetic Variables: = FOS Eclipse of 4 Rosa) The Faint Object Spectrograph ( Blue side) was used to make Rapid = Mode spectrographic observations of the accretion streams in the magnetic = variable star (i.e., Cataclysmic variable) RXJ0515+01 (aka, 4 Rosa). This = is a binary system, and it contains a white dwarf star, an accretion = disk, a mass stream, and hot spot where the mass stream hits the disk and = a companion star. The observations were completed as planned and no = anomalies were reported. 1.5 Completed Two Sets of WF/PC-2 6346 (Snapshot Survey of the = Ultraluminous IRAS Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxies QDOTULIR22 and IR16007+37. These data will be = searched for multiple cores, which imply a merger has occurred, and for = starburst regions. The observations were executed as scheduled, and no = anomalies were noted. 1.6 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red side) was used to make observati= ons of the QSO/radio galaxy 3C395 in order to make a survey of the = Lyman-Alpha line and the 68 Mg II absorption line system. The = observations were completed as scheduled, and no problems were reported. 1.7 Completed FOS/BL/GHRS/2 5959 (Intergalactic Gunn-Peterson = Absorption in Five Quasars) The Faint Object Spectrograph (Blue Side) was used to make = moderate resolution, high signal to noise ratio, spectrographic = observations of the quasar QSO-194025 in order to study its He II = Gunn-Patterson absorption lines. The requested resolution will allow = absorption due to the Lyman-Alpha forest to be separated from absorption = due to the intergalactic medium. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: A guide star loss-of-lock was encountered with FGS #1 with a torque = star selector compensation error type B at 226/0514Z. During the FOS = observation 14 a binary acquisition guide star loss-of-lock occurred = taking the TDF down until the occultation at 0556Z. (See Tomorrow's = report for more details, Item 1.1) The NSSC-1 had a status buffer message at 226/0524Z due to no TDF at = the start of the FOS observations #15. The Shutter did not open and the = observation was lost. The FOS observation #16 was also conducted with the = shutter closed. (See Tomorrow's report for more details, Item 1.1) The ESS took a power hit at 226/0025Z. All ESS systems were back on = line by 226/0300Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH